OpenStudy (babyslapmafro):

Please simplify the following with steps, thank you.. ((x^4)-1)/(x-1)

OpenStudy (anonymous):

The top is a difference of squares \[a^2-b^2=(a+b)*(a-b)\]

OpenStudy (anonymous):

\[\frac{(x^4-1)}{(x-1)}=\frac{ ((x^2)^2-1^2 }{ x-1}=\frac{ (x^2+1)*(x^2-1) }{ x-1 }\]

OpenStudy (anonymous):

then we have a new difference of squares, again in the numerator

OpenStudy (anonymous):

\[x^2-1^2=(x+1)*(x-1)\]

OpenStudy (anonymous):

this makes the equation become \[\frac{(x^2+1)∗(x^2−1^2)}{x−1}=\frac{ (x^2+1)*(x+1)*(x-1) }{ (x-1) }\]

OpenStudy (anonymous):

Last step is the cancel out the same binomials from the top and bottom of the fraction

OpenStudy (babyslapmafro):

(x2+1)∗(x+1)

OpenStudy (babyslapmafro):

Thank you